OpenStudy (anonymous):

a bag contains 4 blue and 3 yellow marbles. the marbles are removed from the bag one at a time without replacement. what is the probability that the fifth marble removed is yellow? express your answer as a common fraction.

OpenStudy (anonymous):

can you please help me...

OpenStudy (raden):

yes i can... :p wait

OpenStudy (raden):

P(B)*P(B)*P(B)*P(B)*P(Y) = 4/7 * 3/6 * 2/5 * 1/4 * 3/3 = ....

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

what is P(B)

OpenStudy (anonymous):

there are other possibilities besides B B B B Y

OpenStudy (raden):

probabilty of event BLUE no, sattelite... the question is the fifth is yellow, so the 1st, 2nd, 3rd, 4th are blue

OpenStudy (anonymous):

@RadEn it didn't say the 5th one is THE FIRST yellow.

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

??

OpenStudy (raden):

there is implicit message, there @r0n4ld

OpenStudy (anonymous):

let y be the number of yellow balls drawn on the first four. clearly y = 0, 1, 2.

OpenStudy (raden):

i got, 24/840 = 1/35 do u have choices

OpenStudy (anonymous):

since there are 4 balls drawn before the 5th which is yellow, the total number of arrangement of y yellow balls and (4-y) blue balls is \[\frac{4!}{y!(4-y)!}\]

OpenStudy (anonymous):

sum these values from y= 0 to 2...: = 11 my answer is \[\large\frac{11}{9\cdot8\cdot7\cdot6\cdot5}\]

OpenStudy (anonymous):

i think that is much snappier than what i had in mind no yellow balls, 5th one is yellow one yellow ball, 5th one is yellow two yellow balls, 5th one is yellow compute these probabilities and add them up is what i had in mind